Showing posts with label card player magazine. Show all posts
Showing posts with label card player magazine. Show all posts

Sunday, April 20, 2014

Poker gems, #470

Ed Miller, in Card Player magazine column, March 19, 2014 (vol. 27, #6), pp. 41-42.


Winning poker requires high aggression and much risk-taking. You cannot systematically reduce the risk and still play a strong game. It can't be done.

Elite players...play a particularly high-variance style because that's the best, most profitable way to play. If you intentionally try to reduce your variance, you must accept that you will never play well.


Monday, February 24, 2014

Poker gems, #469

Andrew Brokos, in Card Player magazine column, January 22, 2014 (vol. 27, #2), page 42.


There's nothing fishier than getting upset about how someone else plays. If you think I'm playing badly, then exploit me and take my money. If I'm giving you tough decisions, then by definition I'm doing something right. Either way, shut up and play the game.

Wednesday, February 12, 2014

Poker gems, #468

Andrew Brokos, in Card Player magazine column, February 5, 2014 (vol. 27, #3), page 46.


If you aren't comfortable backing your reads with the money you have on the table, then you have too much money on the table.

Monday, December 10, 2012

Dealers and players

I very much liked Andrew Brokos's blog post today at CardPlayer.com, musing about player conduct and dealer conduct. It sounds to me as if he behaves at the table much like I do. I've never met him, but I'm pretty sure we would get along famously.

Read the whole thing here.


Saturday, December 08, 2012

How do they know?

The new issue of Card Player magazine (November 28, 2012; vol. 25, #24) has a two-page article by Craig Tapscott on pages 28-29 in which he interviews Alex "Assassinato" Fitzgerald about the latter's thought processes during a hand he played in one of the 2012 WCOOP events.

Here's the second page:



It's the graphic that has me puzzled. If you read the interview, you see that the hand ended when Fitzgerald bet on fourth street and his opponent folded.

So where do they get the river 2c in the graphic? And how do they know that the villain had the Ks-Qs?

My first thought was that the whole graphic was an error, meant for another hand. But no--the cards shown for Fitzgerald perfectly match the text of the article, as do the flop and turn cards.

I also thought that maybe the opponent showed his cards before mucking them, and/or that this occurred on one of the online poker sites that allow rabbit-hunting, in order to see what the last card would have been. But that can't be, because they specify that this occurred during the WCOOP, which is on PokerStars. Stars does not have rabbit-hunting, and does not allow players to show their hole cards when folding.

So who put those false, extraneous pieces of information into the graphic? And why?

Also, am I the only one who notices stuff like this?


ADDENDUM: See comments for explanation of the hole cards, though the 2c river remains a mystery.

Monday, November 19, 2012

Quiz

Steve Zolotow presents an interesting little quiz in his column for Card Player magazine (November 14, 2012, issue; vol. 25, #23, page 50). He asks the reader to estimate what percentage of the time 10s-9s will win against the following hands, assuming all-in before the flop:

1. Ac-Ad
2. Ac-Kc
3. Ac-5c
4. 5c-5d
5. 7c-6c
6. 9c-8c
7. Top 20 percent range
8. Top 40 percent range
9. Top 60 percent range
10. All hands--100 percent range

If this piques your curiosity as it did mine, take a couple of minutes to jot down your answers, before scrolling down.

















Here are the answers Zolotow gives, as percentages, followed in parentheses by the estimates I had written down:

1. 22.8 (20)
2. 38.7 (40)
3. 45.9 (45)
4. 52.4 (50)
5. 62.8 (65)
6. 69.3 (75)
7. 35.4 (30)
8. 41.6 (35)
9. 46.5 (40)
10. 54.0 (45)

Clearly I did a lot better putting in estimates against specific starting hands than I did against ranges. I have a bazillion times more experience with the former than with the latter. For the former, my guesses were all good enough that they would produce the right answer in a go/no-go situation at the table nearly every time. For the latter, though I wasn't outrageously off, I still missed by enough to cause an erroneous decision in some significant spots.

And, frankly, estimating hand strength against an opponent's range is a more important skill than estimating it against a specific hand. It's something I ought to work on.

How about you?

Friday, November 16, 2012

Poker gems, #464

Andrew Brokos, in Card Player magazine column, November 14, 2012 (vol. 25, #23), page 40.


I'd rather have a good hand and the best position than the best hand and poor position.

Monday, October 22, 2012

Poker gems, #463

Ed Miller, in Card Player magazine column, October 17, 2012 (vol. 25, #21), page 36.


If it's nearly impossible against certain players to get your sets paid off for 200 big blinds, doesn't it follow that your monster bluff is a near lock to work against these same players?

Friday, September 14, 2012

I'll take Geography for $500, Alex




Either Card Player magazine and its columnist, Bernard Lee, are confused about what constitutes the "West Coast," or I am. I was not aware that any part of the states of Nevada or Colorado had been given ocean views. Must be a global warming thing.

Friday, September 07, 2012

Card Player takes another crack at it

Apparently somebody at Card Player magazine saw my post from yesterday and took a stab at fixing some of the spelling and punctuation problems I listed in their news story. But while I identified where the problems were, I did not specify what the correct form should be, leaving the staff writers at Card Player to figure this out for themselves. This task proved to be too much for them.


They corrected the two instances of it's where its was called for. They apparently still haven't reset their clocks, and left "PST" (for Pacific Standard Time) unchanged. They corrected getting the run around to getting the runaround.

But the best part, the one that made me literally LOL, was this: They "corrected" back peddled to back-peddled!

Who are these people, and how did they ever get jobs in the writing business?

Thursday, September 06, 2012

Does Card Player magazine have any copy editors?

Card Player's print magazine is pretty bad in terms of copy editing, but the web site is just awful. From just one news story today, I bring you these examples:

The Partouche Poker Tour guaranteed an ambitious €5 million prize pool for it’s [sic] annual main event in Cannes....

Instead of providing the players with an overlay, it appears that tournament officials have back peddled [sic]....

After getting the run around [sic] from the floor, many started playing the tournament under protest....

There were rumors circulating various social media sites that the tour would reluctantly do the right thing and put up the overlay, but as of 9:00 a.m. PST, the prize pool remains unchanged. [Apparently Card Player has not yet gotten around to changing its clocks to daylight savings time.]

By all accounts, it appears the tour has taken it’s [sic] ball and gone home.
Maybe my post title asks the wrong question. Maybe it's too ambitious. Maybe I should instead ask, "Did anybody at Card Player magazine manage to pass fifth-grade English?"

Sunday, August 05, 2012

"Show pool"

Bob Ciaffone is a columnist for Card Player magazine. In his column for the July 25 issue, he is describing a cash game in which he played recently:

About an hour later, I got all-in onA-K again (this time unsuited) in a three-way pot against Q-Q and 2-2. Trip deuces won this pot and I "won" the show pool.
I've never heard this phrase before, and can't figure out its meaning from the context.

Anyone? Anyone? Bueller?


Wednesday, July 25, 2012

Poker gems, #460

Gavin Griffin, in Card Player magazine column, July 11, 2012 (vol. 25, #14), page 60.



Losing players are trying to make hands, winning players are trying to make money.

Tuesday, July 24, 2012

Poker gems, #459

Adam Schoenfeld, in Card Player magazine column, July 11, 2012 (vol. 25, #14), page 52.



No decision in limit hold'em should take more than ten seconds.

Thursday, July 19, 2012

Poker gems, #458

Matt Matros, in Card Player magazine interview, July 11, 2012 (vol. 25, #14), page 34. Deep into a World Series of Poker event earlier this summer, he got all in with K-K versus an opponent's A-K, and lost the massive pot when an ace fell on the river.



There are going to be times when you go from the big stack to the short stack for reasons outside of your control and hopefully you've been through it enough to make a quick mental adjustment, shrug it off and continue to play your best. Having said that, I don't think I've ever experienced anything quite so drastic. Had I won that pot, I would've been the chip leader with about 1 million while second place held only 300,000. Instead, I was left with a below average stack.... I was a little shaken, but I started to look at the positives. I told myself that I was still in the hunt with 30 players left [out of] a massive field. Chances like that don't come around everyday [sic], so I could either sit back and complain about my bad luck, or really try to make the most of the opportunity.


[Matros recovered and went on to win the tournament ($1500 6-max no-limit hold'em), his third bracelet in three years.]

Wednesday, July 11, 2012

Probability of a bigger pair

Ed Miller's column in the June 27 issue of Card Player magazine is about bet-sizing tells. He walks the reader through a hand that he played while holding Q-Q, showing how opponents' bet sizes helped him deduce what hands he was up against. I have no quibbles with his reasoning. In fact, it's a highly worthwhile piece, and the sort of insight that makes me never want to miss his columns.


I do, however, have a quibble with an incidental point tossed in along the way. In explaining why he elected to just call rather than reraise when faced with an opening raise from a nit under the gun, he says:
[T]here are still five unknown hands behind me. There's about a one percent chance each player has either A-A or K-K, making about a five percent chance in total that one of the two hands I'm most afraid of is lurking behind me.
(Note: this is just one of the reasons he cites for not reraising, not the only one.)

At first I wondered why he was making a distinction between the players who have already acted and those yet to act. After all, A-A or K-K is an ugly problem for Q-Q no matter where on the table it is being held. But I think I see what he means. If he were in the big blind, and thus last to act on the raise, and nobody had reraised yet, then he could quite confidently assume that nobody was holding A-A or K-K, unless it is the UTG nit.

Or put it another way: Before there is any action, there is a fixed probability that any other player has been dealt a pocket pair higher than Q-Q. But people so reliably raise and reraise with A-A or K-K that when they fail to do so, we can safely revise downward that estimated probability.

What is the a priori probability that somebody at a ten-handed table has been dealt A-A or K-K when I'm holding Q-Q? It turns out to be a ferociously complicated mathematical problem, one that I could never solve on my own. Fortunately, though, somebody a lot smarter than me has already worked it out, and you can see both the process and the result here:


The table containing the solution is at the bottom of the page. The answer for Q-Q is 0.0841. That is, if I see Q-Q in my hand, and I have nine opponents, about 8.4% of the time one of them will be holding a pair bigger than mine. That does indeed work out to a round estimate of about a 1% probability that any particular one of them has the goods. And that's a useful rule of thumb. Before there is any action (action constituting additional information to throw into our estimation math), each player in a full ring game has about a 1% chance of having been dealt a pair higher than my queens.

If I'm in the big blind and nobody has raised, then probably nobody has a bigger pair, unless they're trying the ol' limp/reraise thing. Also, if I open-raise from under the gun with my Q-Q and get only callers with no reraise, I can go to the flop reasonably confident that nobody has A-A or K-K. But really, that fact has everything to do with what we know from experience about how people play, and little to do with the math. The conclusion would be equally true if the a priori probability of somebody holding a bigger pair were 80%, or 0.8%, rather than the actual 8%.

But I think there's a problem with how Ed applies the math to the situation he describes. Once a nitty player has opened for a raise from under the gun, we no longer have the baseline 8% chance that somebody at the table has A-A or K-K. The probability now has to be estimated to be vastly greater than that. Exactly how much greater depends, obviously, on exactly how nitty this particular nit is. If he is the extreme case and will only raise UTG with exactly A-A or K-K, then our previous 8% probability goes to 100%. Of course, it's not often you'll have a player with a range that narrowly definable, so it usually won't go up all the way to 100%. But it's surely a damn sight higher than 8%.

If the nit does, in fact, have A-A or K-K, then the probability that any other player at the table also has A-A or K-K is much lower than the roughly 1% apiece that we would otherwise be using as our rule of thumb. Using the next column in Alspach's table, we see that the a priori probability that two players have been dealt pocket pairs bigger than my queens is only about 0.2%--a negligible quantity. The implication is this: The fact that a nit put in an UTG raise means that the probability of a bigger pocket pair being out there is much greater than the baseline 8% (as just discussed), but also that nearly all of that probability resides in the nit, not in the players who will be acting behind me.

It is thus no longer true that each player behind us still has a 1% chance of holding aces or kings. Our estimate of the probability that any one of them has A-A or K-K must be greatly reduced from its baseline ~1% by the fact of having observed that UTG raise from the nit. Put most simply: As the probability that the UTG nit has aces or kings goes up, the probability that anybody else at the table has aces or kings goes down.

Now, this is where it gets a little weird and mind-bendy. Because two things are simultaneously true: (A) The probability that any player is dealt A-A is independent of whether any player is dealt K-K.* (B) Alspach's calculation of a 0.2% probability of two players having A-A or K-K is no longer valid once we actually know that one player has one of those hands. The 0.2% is valid only when we have no information except for having seen our own Q-Q.

This is analogous to questions such as the probability of a roulette wheel hitting 00 twice in a row. It is 1/38 x 1/38 = 0.026 x 0.026 = 0.0007. But if we spin and the ball falls in that 00 slot the first time, now the probability of seeing 00 twice in a row greatly increases, to 0.026. Of course, the probability that the second spin will be 00 has not changed one smidgen by virtue of the result of the first spin, yet the probability of seeing two 00s in a row has changed.

So it is with the A-A and K-K. If we know that the UTG player surely or almost surely has A-A or K-K, we have to throw out the window the Alspach table numbers. They no longer apply. The 8% probability (that exactly one player has A-A or K-K when I have Q-Q) is no longer valid because we have strong reason to believe that this specific shuffle of the deck did, in fact, result in one of those 8% events. And the 0.2% probability (that two players have each been dealt A-A or K-K when I have Q-Q) is no longer valid both for that reason, and for the reason that once the UTG nit is known to have A-A or K-K, two of those eight cards are no longer available to be given to other players.

The knowledge (or at least high suspicion) that UTG nit has one of the two premium pairs is akin to opening the sealed box in which Schrodinger's poor cat has been confined. It is now known to be either alive or dead, not in some hybrid or probabilistically indeterminate state.

Mind you, I don't claim to have worked out the math so as to be able to tell you exactly what the new, revised probability is that each of the five players yet to act has a pair bigger than Q-Q, in addition to the one that we presume is hiding under the nit's card protector, but it is definitely not well estimated by the 1% rule of thumb. It must be substantially lower than that. Exactly how much lower depends entirely on the degree of confidence with which we can assign A-A or K-K to the open-raising nit.



*Note: I realize that that is not quite true, because as soon as two aces are in a player's hand and no longer in the deck from which the dealer is pitching cards, the concentration of kings in the deck rises slightly by virtue of card subtraction, so the probability of K-K goes up slightly. But it's a tiny difference.

Sunday, June 24, 2012

When to look at your cards?

Steve Zolotow, in his column for Card Player magazine, June 13, 2012, writes this (page 40):

I hope everyone avoids looking at their cards until it is their turn to act. There are two reasons for this. First, you can't reveal anything about your hand because you don't know it yet. Second, you can focus on your opponents and not on your cards.
I have, of course, heard this argument before--many times, in fact. When I was first learning about poker by watching television shows like World Poker Tour and Celebrity Poker Showdown, I specifically remember hearing both Phil Gordon and Annie Duke explain that looking at one's cards before it is one's turn is an amateurish mistake.

I disagree.

I understand perfectly well the arguments made in favor of waiting to look. But what advocates of this position usually omit, or perhaps do not grasp, is that there are countervailing arguments.

First, waiting to look until it's your turn means that the whole game runs more slowly. That is not a small consideration.

Second, to quote Phil Ivey, "If you have tells, work on them." If you're giving away information by body language or facial expressions when you check your hole cards, that's a problem you need to fix pronto. Whether you give off that information a few seconds after you get your cards or only when it's your turn makes little difference. You need to not be spewing such tells, period. Of course it is the case that your action (fold, call, or raise) will also be giving off information, but that is a separate matter. You don't want your face and body to be saying one thing and your bet a contradictory thing.

Third, if you wait to look until it's your turn, all eyes will be on you as you do so, because everybody is waiting for you to act. Thus, if you do have tells, the way to be sure they get maximal attention from opponents is to wait until everybody is watching you while you check your hole cards. On the other hand, if you do it when most other players are doing it--as they receive their last down card--they're all busy with their eyes elsewhere.

Fourth, if you wait until it's your turn, you're likely to give off information by how long it takes you to decide what to do after looking at your cards. Raise or not? If so, how much? Even if your face reveals nothing, the delay in action can suggest how complicated or difficult your decision is. I say let that decisional time pass while other things are going on.

I don't usually like appeals to authority in argument, but it is definitely not the case that all the best players in the world wait to look until it is their turn to act. These days we are getting more and more live broadcasts of late stages of big poker tournaments. Just watch--you'll see that there is a lot of variation on this point among the greats. Though T.J. Cloutier seems to be pretty much vanishing from the tournament scene now that he is, sadly, in the twilight of his life, I clearly remember him addressing this controversy in one of his books. He said that he likes to give a quick glance at his cards as soon as they arrive, then direct his attention at what how the other players are reacting to theirs. I agree with him. It takes literally one second or less, so you won't be missing much.

I like having my own cards in mind as I watch the other players act. Often my default decision about what to do changes. A completely crap hand becomes raiseworthy if I'm in late position and everybody has folded. Marginal hands like A-8 and Q-J can go from being planned raises to being folds if there is any action before it gets to me, or reraises if a raise comes from a habitually frisky player. And so forth. I would be stunned beyond belief if any of my opponents can read on my face these changes in my thinking as the action unfolds.

For what it's worth, I've noticed that Cardgrrl adopts a compromise position. She looks at her cards when the action is about two seats to her right. There is an argument to be made that this tactic incorporates some of the advantages of both of the more common approaches (looking immediately, or not until it's your turn). It's worth considering.

I don't mind that there are differences of opinion, preference, and style on this question. What bugs me, though, is how dogmatic many proponents of the Zolotow/Gordon/Duke position are about it--insisting that theirs is the only right way and any deviation from it is a clear error. I just don't buy it.

Saturday, June 09, 2012

Steve Zolotow is a fool!

Steve Zolotow's column in the May 30 issue of Card Player magazine is about becoming a tricky player, one who is difficult to read. His point is that you have to play different starting hands the same way as each other, and also play the same starting hand different ways at different times.


For his first example, he takes A-A and discusses the default play (reraise) plus some alternatives. He then does the same for suited K-J, medium suited connectors, and a small pair--in each case mentioning the default play plus how they should sometimes be played as if the cards are something else, like A-A.

Fine so far.

But then he screws up. For his last example, he uses 4s-2h. Yes, the Mighty Deuce-Four. Of this he says, "With 4-2 offsuit, your default play is to fold. But occasionally you can raise as a pure squeeze play, hoping everyone folds. If you get re-raised, it is an easy fold."

The man does not know whereof he speaks. He is clearly unqualified to be writing a poker column. I demand that he be fired at once!

However, he does mention something I had not heard before, which he says is "an old pun": When one wins with the 4-2, one can say, "It was a fortuitous hand."

Get it?

Four-two-itous.

Friday, May 25, 2012

Poker gems, #456

Ed Miller, in Card Player magazine column, May 16, 2012 (vol. 25, #10), page 38.



Whenever you beat players in a hand and they complain about how you played it, you're officially in their head.

Poker gems, #455

Andrew Brokos, in Card Player magazine column, May 16, 2012 (vol. 25, #10), page 37.



There is one reason to check and call on the river, and that's to induce bluffs. If you don't expect your opponent to bluff the river, then your options are either to value bet the hand yourself or to check and fold. Checking and calling makes no sense in this situation because virtually any player's calling range will be wider than his value betting range.